# Reconstructing Satellites in 3D from Amateur Telescope Images

### v3

@@ 4.1.4 Joint 3D Reconstruction and Pose Refinement
We then compare our pipeline with two NeRF-based baseline methods, NeRF--- - and BARF, that support joint 3D reconstruction and pose estimation. To ensure a fair comparison, we use our customized pose initialization for all baselines; otherwise, neither NeRF--- - nor BARF exhibits generalization ability. In addition, traditional NeRF-based methods uniformly sample the space between the cameras and the object (illustrated by the light blue dots in Fig.6). This approach fails in our sparse-view, long-distance observational scenario because of overfitting in the excessively large sampling space. Consequently, we explore a BARF variant that restricts the sampling region to the vicinity of the initial object distance predicted by our modified SfM (indicated by the gray dots in Fig.6). This adaptive strategy significantly reduces overfitting and improves reconstruction accuracy.
Figure 7 compares the reconstructed point clouds and novel views from our method with those produced by baseline techniques. Although NeRF-based methods have been carefully tuned for this task, they still struggle to generalize to unseen viewpoints. In contrast, our GS-based approach employs controlled point growth to focus sampling within the object region, accurately capturing fine structural details. This advantage is confirmed by superior SSIM, PSNR, and LPIPS metrics (Table III and Fig.5). In the meantime, our method delivers more accurate pose estimates after the joint optimization of poses and the 3D object (Table II).
@@
Our method also significantly reduces training time. A 3D satellite model is reconstructed in approximately 4 minutes on a single NVIDIA RTX 4090 using our method (2 minutes for GS training and 2 minutes for BnB pose refinement), much faster than NeRF--- -’s 1.5 hours and BARF’s over 4 hours.
@@ 4.2 On-Sky Results
Figure 9: An official schematic of the Chinese Space Station (CSS). The CSS comprises four primary modules: 1 - Tianhe Core Module, 2 - Mengtian Laboratory Module, 3 - Wentian Laboratory Module, 4 - Tianzhou-6 Cargo Spacecraft.
4.2.1 Observations of CSS and ISS
The CSS was observed on September 15, 2023, from 20:48:30 to 20:51:17 UTC near Miyun Reservoir, Beijing, with a closest approach of 389.4 km and a peak elevation of 78.3°. The ISS was observed on August 21, 2024, from 20:20:06 to 20:22:09 UTC in Yanqing District, Beijing, with a closest approach of 442 km and a peak elevation of 86°. For both observations, we used a Celestron C14HD telescope (0.35 m diameter, f/11) paired with a QHY5III678M camera (Fig.11), achieving a spatial sampling rate of 0.106 arcsec/pixel (approximately 0.2 m/pixel). Our observations employed the ”Fake Polar Axis Method” with a German equatorial mount (CGX-L), uniquely oriented to a precisely calculated azimuth relative to the satellite’s track rather than true north. This innovative approach eliminates the need for meridian flips and minimizes field rotation, particularly in the center of the image sequence. Images were captured at 5 ms exposures, averaging 88 fps for CSS and 55 fps for ISS, resulting in datasets of 13,800 images for CSS and 6,770 for ISS, with resolutions of 360 and 640, respectively.
(a) 3D Point Cloud Reconstruction Results
(a) GT
(b) Ours
(c) Original GS
(d) No Preprocessing
(e) No Regulation
(f) No BnB
(g) No Filtering
(b) Novel View Synthesis results
(h) GT
(i) Ours
(j) Original GS
(k) No Preprocessing
(l) No Regulation
(m) No BnB
(n) No Filtering
@@ 4.2.1 Observations of CSS and ISS
Figure 11: The imaging instrument. The Celestron C14HD telescope paired with a QHY5III678M planetary camera
@@
We first apply the pre-processing pipeline to extract clean frames from the raw observations, selecting 138 clean frames for CSS and 116 for ISS. Following our simulation protocol, we chose one image out of every ten for training—yielding 14 images for CSS and 12 for ISS—with the remainder reserved for validation. Our method is again benchmarked against three NeRF-based baselines, NeRF--- -, BARF and BARF’s variant (BARF with Constrained Sampling, abbreviated as BARF (Con.)).
@@ 5 Conclusion
Due to hardware limitations, such as the telescope’s aperture size, our current 3D imaging resolution remains restricted. In future work, we plan to use larger, more stable telescopes to improve image resolution and signal-to-noise ratio, thereby enabling the reconstruction of smaller satellites. We also aim to incorporate material properties into our 3D satellite modeling to more accurately simulate sunlight reflection and occlusion effects. Additionally, we plan to extend our method to other 3D imaging applications characterized by long observation distances and low image quality.

### v5

@@ 4.1.4 Joint 3D Reconstruction and Pose Refinement
We then compare our pipeline with two NeRF-based baseline methods, NeRF-- and BARF, that support joint 3D reconstruction and pose estimation. To ensure a fair comparison, we use our customized pose initialization for all baselines; otherwise, neither NeRF-- nor BARF exhibits generalization ability. In addition, traditional NeRF-based methods uniformly sample the space between the cameras and the object (illustrated by the light blue dots in Fig. 8). This approach fails in our sparse-view, long-distance observational scenario because of overfitting in the excessively large sampling space. Consequently, we explore a BARF variant (named BARF (Con.)) that restricts the sampling region to the vicinity of the initial object distance predicted by our modified SfM (indicated by the gray dots in Fig. 8). This adaptive strategy significantly reduces overfitting and improves reconstruction accuracy.
Fig. 9 compares the reconstructed point clouds (appearance and shape attributes of 3D Gaussians neglected) and novel views from our method with those produced by baseline techniques. Although NeRF-based methods have been carefully tuned for this task, they still struggle to generalize to unseen viewpoints. In contrast, our GS-based approach employs controlled point growth to focus sampling within the object region, accurately capturing fine structural details. This advantage is confirmed by superior SSIM, PSNR, and LPIPS metrics (Table III and Fig. 6). In the meantime, our method delivers more accurate pose estimates after the joint optimization of poses and the 3D object (Table II).
@@
Our method also significantly reduces training time. A 3D satellite model is reconstructed in approximately 4 minutes on a single NVIDIA RTX 4090 using our method (2 minutes for GS training and 2 minutes for BnB pose refinement), much faster than NeRF--’s 1.5 hours and BARF’s over 4 hours.
Figure 11: The Chinese Space Station (CSS) and the imaging instrument. On the left, the official schematic of the CSS, featuring four primary modules: 1 - Tianhe Core Module, 2 - Mengtian Laboratory Module, 3 - Wentian Laboratory Module, and 4 - Tianzhou-6 Cargo Spacecraft. In the lower right corner, the Celestron C14HD telescope paired with a QHY5III678M planetary camera.
@@ 4.2 On-Sky Results
(a) 3D Point Cloud Reconstruction Results
(a) GT
(b) Ours
(c) Original GS
(d) No Preprocessing
(e) No Regulation
(f) No BnB
(g) No Filtering
(b) Novel View Synthesis results
(h) GT
(i) Ours
(j) Original GS
(k) No Preprocessing
(l) No Regulation
(m) No BnB
(n) No Filtering
@@
4.2.1 Observations of CSS and ISS
The CSS was observed on September 15, 2023, from 20:48:30 to 20:51:17 UTC near Miyun Reservoir, Beijing, with a closest approach of 389.4 km and a peak elevation of 78.3°. The ISS was observed on August 21, 2024, from 20:20:06 to 20:22:09 UTC in Yanqing District, Beijing, with a closest approach of 442 km and a peak elevation of 86°. For both observations, we used a Celestron C14HD telescope (35cm aperture, f/11) with a QHY5III678M camera (shown in the lower right corner of Fig. 11), achieving a spatial sampling rate of 0.106”/pixel (0.2m/pixel at 400km). This sampling rate is smaller than the optical resolution: The diffraction limit of our 35 cm telescope is approximately 0.39”, corresponding to 76cm resolution at 400km. Our observations employed the ”Fake Polar Axis” method with a German equatorial mount (CGX-L), whose polar axis was tilted into the satellite’s orbital plane rather than aligned with the celestial pole (see Fig. 7 (b)). This orientation eliminates the need for meridian flips and enables uninterrupted tracking during the satellite pass. As most of the satellite’s apparent motion is projected onto the declination (Dec) axis, only minor compensation is required from the right ascension (RA) axis, reducing field rotation—especially near the center of the image sequence. We implemented this method using a custom tracking program. Images were captured at 5 ms exposures, averaging 88 fps for CSS and 55 fps for ISS, resulting in datasets of 13,800 images for CSS and 6,770 for ISS, respectively. The relative angular sweeps between the telescope and the satellites span approximately 99° for CSS and 82° for ISS.
@@ 4.2.1 Observations of CSS and ISS
We first apply the pre-processing pipeline to extract clean frames from the raw observations, selecting 138 clean frames for CSS and 116 for ISS. Following our simulation protocol, we chose one image out of every ten for training—yielding 14 images for CSS and 12 for ISS—with the remainder reserved for validation. Our method is again benchmarked against three NeRF-based baselines, NeRF--, BARF and BARF’s variant (BARF with Constrained Sampling, abbreviated as BARF (Con.)).
@@ 5 Conclusion
Due to hardware limitations, such as the telescope’s aperture size, our current 3D imaging resolution remains restricted. Additionally, our current model, based on a simple photometric-consistency loss, does not fully capture strong specular reflections from metallic surfaces and solar panels, which may bias the fit under highly specular conditions. In future work, we plan to use larger, more stable telescopes to improve image resolution and signal-to-noise ratio, enabling the reconstruction of smaller satellites. We will also develop modular interfaces so that updated denoising and matching networks can be seamlessly integrated. Additionally, we plan to incorporate material properties into our 3D satellite models to better simulate sunlight reflection and occlusion effects. We will further extend our approach to capture specular reflections by using reflectance-aware Gaussian splatting. Additionally, we will explore other applications of our method for 3D imaging at long observation distances and with low-quality images.
Acknowledgments
This work was supported by National Natural Science Foundation of China (Grant No. 62371007, 62136001, 62088102, 12173041).